solve this system of equations by elimination
3x+4y=13
5x-9y=6

Answer by ly12603(44)

y=1 x=3 start by solving for one of the variables:
3x+4y=13
-4y+-4y
3x=13-4y
(divide by 3)
x=4.3(repeating)-1.3(repeating)y
*next substitute what you just found out in*
5x-9y=6
5(4.3(repeating)-1.3(repeating) y)-9y=6
21.6 (repeating)-6.6 (repeating)y-9y=6
-21.6(repeating)=-21.6 (repeating)
*equals*
-15.6(repeating) y = -15.6 (repeating)
**divide by the coefecient of y**
y=1
*next substitute what you found out for y in*
5x-9y=6
5x-9(1)=6
+9=+9
5x=15
*divide by 5*
x=3